# Urban Region Pre-training and Prompting: A Graph-based Approach

### v1

@@ Abstract.
Urban region representation is crucial for various urban downstream tasks. However, despite the proliferation of methods and their success, acquiring general urban region knowledge and adapting to different tasks remains challenging. Previous work often neglects the spatial structures and functional layouts between entities, limiting their ability to capture transferable knowledge across regions. Further, these methods struggle to adapt effectively to specific downstream tasks, as they do not adequately address the unique features and relationships required for different downstream tasks. In this paper, we propose a Graph-based Urban Region Pre-training and Prompting framework (GURPP) for region representation learning. Specifically, we first construct an urban region graph that integrates detailed spatial entity data for more effective urban region representation. Then, we develop a subgraph-centric urban region pre-training model to capture the heterogeneous and transferable patterns of interactions among entities. To further enhance the adaptability of these embeddings to different tasks, we design two graph-based prompting methods to incorporate explicit/hidden task knowledge. Extensive experiments on various urban region prediction tasks and different cities demonstrate the superior performance of our GURPP framework. The implementation is available at this repository: https://anonymous.4open.science/r/GURPP.
Urban Computing, Urban Region Representation Learning
copyright: acmlicensed
@@ 1. Introduction
An urban region is a geographical area comprising diverse spatial entities (e.g.,  shopping centers, restaurants, and roads) and their interactions (e.g.,  nearby and located in). Recently, urban region representation learning, which aims to extract vector embeddings from these entities, has become a significant research focus (Zhang et al., 2024). Effective urban region representations can be applied to various tasks, such as region popularity prediction (Li et al., 2024; Fu et al., 2019; Zhou et al., 2023a; Wu et al., 2022), house price prediction (Wang and Li, 2017), crime prediction (Wang and Li, 2017; Zhou et al., 2023a; Wu et al., 2022; Zhang et al., 2021), and population density inference (Li et al., 2023), enabling urban planners, policymakers, and businesses to enhance decision-making processes. However, the complexity of urban environments and the diversity of applications pose significant challenges.
@@
Balancing urban generality and task specificity is crucial in urban region representation. (i) Urban regions often appear similar in their spatial structures and functional layouts, because the interactions among entities within these regions follow similar patterns. These shared patterns serve as generic knowledge across urban areas, facilitating the development of generalizable region representation models. As illustrated in Figure 1, Region A and Region B showcase typical urban regions. In these examples, spatial structures are demonstrated by the proximity of residences to parks and shops, which enhances community convenience and accessibility. Functional layouts are characterized by the clustering of similar entities within areas, such as shops in commercial areas and schools in teaching areas, and the close proximity of these distinct areas to each other. Therefore, an effective urban region representation model needs to capture the heterogeneous and generic patterns of interactions among entities within regions. (ii) Despite the broad similarities, each urban region and downstream task possesses unique details. Specific tasks may require attention to different features within urban data. For example, in traffic crash prediction, the density of roads and the number of residents living nearby, might be critical features, as they directly influence the probability of accidents occurring. On the other hand, crime prediction models might focus on the density and type of POIs within the region, such as bars or late-night convenience stores. As consequence, balance between generality and specificity raises a crucial question: Can we design an urban region representation framework that effectively integrates generalizable knowledge while accommodating the specific needs of distinct tasks?
Recently, the pre-training and prompting paradigm that learns general representations that can be fine-tuned or prompted to adapt to specific tasks has shown remarkable success in the field of natural language processing (Liu et al., 2023b; Brown et al., 2020; Roberts et al., 2020; Gao et al., 2021) and computer vision (Radford et al., 2021; Zhou et al., 2022b, a). Inspired by these advancements, we hope to design a pre-training and prompting framework for urban region representation learning. However, building such a framework for urban representation is never easy. The challenges are as follows.
(1) Capturing the heterogeneous and generic patterns of interactions among spatial entities within regions is hard. Most previous methods follow a process where they first predefine aggregations and statistics about urban entities and relationships, and then learn from these predefined features (Han et al., 2020; Yang et al., 2022; Zhang et al., 2023; Li et al., 2023). While this approach has its merits, it often fails to capture the spatial structures and functional layouts between entities within regions which is generalizable. Other previous methods consider urban entities in isolation, but they focus on only a few types, such as POIs and roads and a few relations between them. This narrow focus limits their ability to capture the heterogeneous and generic patterns of interactions among entities, resulting in embeddings that are less robust and less effective (Fu et al., 2019).
(2) Tuning pre-trained region embeddings to capture the tasks’ focus on different features or relationships within the original urban data poses a challenge. The majority of previous urban region representation methods use simple regression models which merely projects the pre-train embeddings without capturing the specific entities and relationships relevant to individual tasks (Li et al., 2023; Zhang et al., 2023; Li et al., 2024; Yao et al., 2018; Wu et al., 2022; Jenkins et al., 2019; Zhang et al., 2021). In contrast, prompting offers a more flexible approach by allowing models to focus on specific aspects of the data as needed (Liu et al., 2023b; Yoo et al., 2023; Sun et al., 2023), effectively highlighting small but crucial differences between urban downstream tasks. Despite this potential, existing prompting methods (Yan et al., 2024; Zhou et al., 2023a) have limitations. These methods use simple prompts that do not consider detailed and heterogeneous interactions among various urban entities. Additionally, many existing prompting techniques simply concatenate prompts onto the original embeddings, lacking the flexibility to selectively filter out entities and relationships that are specifically useful for downstream tasks. This limitation makes it challenging to tailor the generalizable embedding precisely to the unique characteristics of different urban downstream tasks.
Figure 2. Main idea of urban pre-training and prompting
In light of the challenges discussed, we propose a novel framework named GURPP (Graph-based Urban Region Pre-training and Prompting Framework) as shown in Figure 2. Specifically:
To address the first challenge, we develop the Subgraph-centric Urban Region Pre-training (GURP) model. To capture the spatial structures and functional layouts between entities, we construct an urban region graph that integrates urban entities and their relationships. Unlike previous methods where each region is treated as a single point, our approach views each region as a subgraph. So the GURP model then extract region subgraphs from urban region graph with graph patterns and learns region embeddings from these subgraphs through a heterogeneous graph learning model. To further enhance the learning process, we design a learning pipeline that incorporates heterogeneous graph learning model and multi-view learning. It enables the model to learn both structure information within subgraphs and the relationships between subgraphs.
To address the second challenge, we develop a knowledge-enriched prompting learning model. This model includes two approaches: a manually-designed prompt method and a task-learnable prompt learning method. The manually-designed prompt adjusts subgraph structures to highlight task-specific relationships and entities using predefined weights. The task-learnable prompt learns from task data to generate prompt graphs that capture implicit task-specific knowledge. By incorporating both predefined and learned task-specific prompts, the model tailors the general embeddings to better suit the unique characteristics and requirements of various urban computing tasks.
Contributions:
  • (Urban region graph). We introduce urban region graph, which integrates detailed spatial and entity data for more effective urban region representation.
  • (Pre-training model). We propose a Subgraph-centric Urban Region Pre-training (GURP) model for learning embeddings from urban region graphs, capturing generalizable patterns of interactions among entities within regions.
  • (Prompting model). We design a novel knowledge-enriched prompting learning model with two graph-based prompting methods: a manually-defined prompt to incorporate explicit task knowledge and a task-learnable prompt to discover hidden task-relevant patterns.
@@
  • (Extensive experiments). We demonstrate the effectiveness of our GURPP framework through extensive experiments on three urban region prediction tasks across two cities, achieving significant performance improvements with an average increase of $17.28\%$ on MAE and $22.26\%$ on RMSE.
2. Preliminaries
@@
Once the mapping function $\zeta(\cdot)$ has been learned, it can be applied to various downstream prediction tasks, such as crime and check-in predictions. Let the task dataset be $\{(r_{i},y_{i})\}_{i=1}^{|\mathcal{R}|}$, where $y_{i}\in\mathbb{R}$ is a numerical indicator for region $r_{i}$. The downstream prediction tasks aim to predict $y_{i}$ based on the region embedding $\zeta(r_{i})$ using a simple regression model like a Ridge Regression (Hoerl and Kennard, 1970).
Figure 3. The Overview of our GURPP. (1) Create urban region graph that integrates specific urban entities. (2) Extract subgraph of each region according to the graph pattern. (3) Use multi-view learning to learn the subgraph representations with GURP model. (4) Adapt the region embeddings with manually-designed prompt method and task-learnable prompt learning method.
3. Overview
As shown in Figure 3, we propose a graph-based framework to pre-train and prompt urban region representations. Our approach begins with the creation of an urban region graph that integrates specific urban entities, providing a comprehensive structure for representing spatial and functional relationships within urban areas (discussed in Section 4.1). From this urban region graph, we extract subgraphs corresponding to each region, using graph patterns that characterize the region’s structure and interactions. These subgraph representations are then refined through multi-view learning using the GURP model, enabling the capture of generalizable information that can serve as the foundation for downstream tasks (explained in Section 4.2). To tailor these representations for specific tasks, we employ two prompt-tuning methods: a manually-designed prompt method and a task-learnable prompt learning method, each designed to incorporate task-specific knowledge into the pre-trained region embeddings (elaborated in Sections 5.1 and 5.2, respectively).
@@ 4.1. Urban Region Graph Construction
An urban region graph is constructed to model the spatial structures and functional layouts of urban areas. The urban region graph, denoted as $G=(V,E,T_{V},T_{E})$, is a heterogeneous graph containing multiple types of nodes and relationships. Here, $v\in V$ and $e\in E$ represent a node and an edge, respectively, and $\phi(v)\in T_{V}$ $(\psi(e)\in T_{E}$) is a type function that assign a type to a specific node (edge).
The urban region graph follows a schema that defines the relationships among various types of nodes and edges (see Figure 4). Without loss of generality, we consider three types of spatial entities: POIs, roads, and junctions. Each spatial entity is categorized: for instance, POI categories include shops and parks, road categories encompass highways and residential streets, and junction categories cover features like roundabouts.
Figure 4. Schema of urban region graph.
We construct the urban region graph from POI datasets and online maps. Since the urban region graph contains heterogeneous nodes and edges, we adopt the knowledge graph embedding technique TransR (Lin et al., 2015) to initialize the representations of all nodes and edges. This aligns the heterogeneous information into a unified vector space.
@@ 4.2. Subgraph-centric Region Embedding
The region-induced subgraph is a subgraph of $G$ that includes the region node $r$ and matches the graph pattern $\mathcal{P}=(P_{V},P_{E},\theta)$. Here, $P_{V}\subseteq T_{V}$ and $P_{E}\subseteq T_{E}$ are sets of node and edge types, respectively, which are used to filter entities and relationships specifically useful for region embeddings. The set $\theta\subseteq T_{V}$ indicates the node types that serve as termination points for subgraph extraction. As a result, the schema of the region-induced subgraph is a subset of the schema of the urban region graph. During the pre-training phase, we select all node and edge types and set $P_{V}=T_{V},P_{E}=T_{E}$, i.e., $\mathcal{P}=(T_{V},T_{E},\theta)$ with the aim of comprehensively characterising the rich information of a region and thereby learning general knowledge.
The region-induced subgraph $g_{r}=(V_{r},E_{r},P_{V},P_{E})$ includes nodes $V_{r}$ and edges $E_{r}$ that can be reached from $r$ and whose types match $P_{V}$ and $P_{E}$. We extract $g_{r}$ recursively, starting with $V_{r}=\{r\}$ and $E_{r}=\varnothing$ and repeatedly expanding $V_{r}$ and $E_{r}$ until the termination points are reached. Specifically, we use a breadth-first search (BFS) starting from $r$ to expand the subgraph. During each iteration of BFS, we explore edges $e(u,v)$ where one endpoint $u$ is in $V_{r}$ and the other endpoint $v$ is not yet in $V_{r}$. We then check whether the type of $e$ is in $P_{E}$ and whether the type of the unseen node $v$ is not in the termination type set $\theta$. If $e$ passes the checks, we add it to the set $\Delta E_{r}$. Using $\Delta E_{r}$, we construct the node set $\Delta V_{r}$, which contains the endpoints in $\Delta E_{r}$ that are not already in $V_{r}$. Formally, the sets of incremental edges (i.e.,  $\Delta E_{r}$) and nodes (i.e.,  $\Delta V_{r}$) are as follows: $\Delta E_{r}=\{e(u,v)~{}|~{}u\in V_{r}\land\phi(v)\notin\theta\land e\in E\land\psi(e)\in P_{E}\}$ and $\Delta V_{r}=\{v|~{}v\notin V_{r}\land\exists u\in V_{r},e(u,v)\in\Delta E_{r}\}$. We then update the region-induced subgraph by setting $V_{r}=V_{r}\cup\Delta V_{r}$ and $E_{r}=E_{r}\cup\Delta E_{r}$. The expansion terminates if $\Delta E_{r}$ and $\Delta V_{r}$ are empty.
@@
In Section 4.1, each node in the subgraph $g_{r}$ has been assigned a global initial embedding using TransR. Here, we encode each subgraph $g_{r}$ into a vector $\vec{h}_{r}$ to further incorporate the local features of region $r$. Specifically, we use a heterogeneous graph encoder $\textsf{gEnc}(\cdot)$ to extract the heterogeneous and generic patterns within $g_{r}$, ultimately outputting the contextualized embedding of the subgraph.
Our encoder $\textsf{gEnc}(\cdot)$ chooses the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as the subgraph node encoder due to HGT’s capability of capturing node and edge-types semantics. By using HGT, $\textsf{gEnc}(\cdot)$ characterizes the heterogeneity of input graph by maintaining a set of node and edge-type dependent parameters, and the output embedding of each node $\vec{v}^{HGT}$ that integrates the global initial embeddings generated by TransR, and node and edge-type semantics of the subgraph $g_{r}$. This trait aligns with our need to capture the semantic features of heterogeneous entities and relationships within each region-induced subgraph.
After obtaining the representations of each node in the subgraph, we further obtain the graph-level representation of $g_{r}$ through two steps of aggregation. For each subgraph $g_{r}$, we first use $\text{SUM}(\cdot)$ function to aggregate the representations of nodes of the same type to integrate the impact of the number of entities on the representation, obtaining the node-type embedding $\vec{\tau}_{i}$:
$\vec{\tau}_{i}=\sum_{\phi(v)=\tau_{i}}\vec{v}^{HGT},\quad\tau_{i}\in T_{V},\forall v\in V_{r}.$
Then we concatenate all representations of node types, and encode them with a Linear layer to capture the interaction between various types of nodes:
$\vec{h}_{r}=\text{Linear}\left(\text{CONCAT}_{|T_{V}|}(\vec{\tau}_{i})\right).$
The output $\vec{h}_{r}$ of $\textsf{gEnc}(\cdot)$ is the final embedding of region-induced subgraph $g_{r}$. Once the graph encoder is pretrained, any graph that conforms to the pattern $\mathcal{P}$ can be encoded by $\textsf{gEnc}(\cdot)$ to generate the corresponding representations, demonstrating high flexibility.
4.3. Multi-view Representation Learning
We learn the region representation with multiple views to characterize general region features by using the spatial proximity, region satellite imagery and region-wise traffic information.

### v4

@@ Abstract.
Urban region representation is crucial for various urban downstream tasks. However, despite the proliferation of methods and their success, acquiring general urban region knowledge and adapting to different tasks remains challenging. Existing work pays limited attention to the fine-grained functional layout semantics in urban regions, limiting their ability to capture transferable knowledge across regions. Further, inadequate handling of the unique features and relationships required for different downstream tasks may also hinder effective task adaptation. In this paper, we propose a Graph-based Urban Region Pre-training and Prompting framework (GURPP) for region representation learning. Specifically, we first construct an urban region graph and develop a subgraph-centric urban region pre-training model to capture the heterogeneous and transferable patterns of entity interactions. This model pre-trains knowledge-rich region embeddings using contrastive learning and multi-view learning methods. To further refine these representations, we design two graph-based prompting methods: a manually-defined prompt to incorporate explicit task knowledge and a task-learnable prompt to discover hidden knowledge, which enhances the adaptability of these embeddings to different tasks. Extensive experiments on various urban region prediction tasks and different cities demonstrate the superior performance of our framework.
Urban Foundation Model, Region Representation, Graph Prompt
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadadoi: XX.XXXX/XXXXXXX.XXXXXXXisbn: XXX-X-XXXX-XXXX-X/XXXX/XXccs: Information systems Data miningccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
An urban region is a geographical area comprising diverse spatial entities (e.g.,  shopping centers and roads) and their interactions (e.g.,  nearby and located in). Recently, urban region representation learning, which aims to extract vector embeddings from these entities, has become a significant research focus (Zhang et al., 2024). Effective urban region representations can be applied to various tasks, such as region popularity prediction (Li et al., 2024a; Fu et al., 2019; Zhou et al., 2023a; Wu et al., 2022), house price prediction (Wang and Li, 2017), and crime prediction (Wang and Li, 2017; Zhou et al., 2023a; Wu et al., 2022; Zhang et al., 2021), enabling urban planners and businesses to enhance decision-making processes.
@@
However, the complexity of urban environments and the diversity of applications pose significant challenges. On the one hand, urban regions often appear similar in their spatial structures and functional layouts, because the interactions among entities within these regions follow similar patterns. These shared patterns serve as generic knowledge across urban areas, facilitating the development of generalizable region representation models. As illustrated in Figure 1, Regions A and B showcase typical urban regions. In these examples, spatial structures are demonstrated by the proximity of residences to parks and shops, which enhances community convenience and accessibility. Functional layouts are characterized by the clustering of similar entities within areas, such as shops in commercial areas and schools in teaching areas, and the close proximity of these distinct areas to each other. Therefore, an effective urban region representation model needs to capture the heterogeneous and generic patterns of interactions among entities within regions. On the other hand, despite the broad similarities, each urban region and downstream task possesses unique details. Specific tasks may require attention to different features within urban data. For example, in traffic crash prediction, the density of roads and the number of residents living nearby might be critical features, as they directly influence the probability of accidents occurring. Meanwhile, crime prediction models might focus on the density and type of POIs (Point-of-Interests) within the region, such as bars or late-night convenience stores. As a consequence, the balance between generality and specificity raises a crucial question: Can we design an urban region representation framework that effectively integrates generalizable knowledge while accommodating the specific needs of distinct tasks?
To solve this problem, this paper presents a simple yet effective framework with pre-training and prompting on geographical data, since this new paradigm has shown remarkable success in balancing embedding generality and task specificity in the areas like natural language processing (Liu et al., 2023c; Roberts et al., 2020), computer vision (Zhou et al., 2022b, a), and graphs (Sun et al., 2023, 2024b). The intrinsic reason behind this superiority, as many researchers reported (Wang et al., 2025), lies in the powerful capability of data manipulation from the prompt side, and knowledge learning from the pre-training side, which can offer flexible adaptation for various downstream tasks and useful knowledge for generalizable representation. However, handling such complex non-linear geographical data requires prior experience. Challenges are twofold:
  • $\circ$
    Difficulties in capturing the heterogeneous and generic knowledge from complex spatial entity interactions. Existing methods mostly predefine aggregations and statistics of some urban entities and relationships, and then use these these predefined features for downstream tasks (Yang et al., 2022; Zhang et al., 2023; Li et al., 2023), which often fail to capture the spatial structures and functional layouts between various entities within regions.
  • $\circ$
    Difficulties in capturing task-specific patterns. Existing methods usually use simple prompts like directly concatenating them to the pre-trained embeddings (Yan et al., 2024; Zhou et al., 2023a), which might work for certain domains but might fall short for urban region tasks with heterogeneous and complex geographical entity interactions.
Figure 2. Main idea of urban pre-training and prompting.
In response to these challenges, we propose GURPP, a Graph-based Urban Region Pre-training and Prompting Framework as shown in Figure 2. Specifically, GURPP leverages the expressive power of urban graphs to capture both generic and task-specific urban features. By constructing the urban region graph, GURPP effectively extracts the common spatial structures and functional layouts inherent to urban areas. Simultaneously, GURPP addresses the specificity required for different downstream tasks through a two-stage process. First, a subgraph-centric pre-training model, enhanced by multi-view self-supervised learning, captures diverse urban patterns—ranging from mobility trends and region imagery to intricate spatial and functional relationships. Then, two complementary graph-based prompting strategies further refine these representations: a manually-designed prompt that incorporates explicit task-specific semantics by emphasizing critical relationships and entities, and a task-learnable prompt that adapts to the hidden complexities of urban data. This design aims to bridge the gap between general urban knowledge and task-specific requirements for urban region representation. Contributions are as follows.
  • $\circ$
    Graph-based Urban Region Framework: We introduce GURPP, a novel graph-based pre-training & prompting framework for urban region representation, which unifies the modeling of complex spatial entity semantics. This design enables the extraction of heterogeneous and generic urban features, overcoming the limitations of methods that rely on predefined aggregations.
  • $\circ$
    Subgraph-centric Pre-training and Dual Prompting Models: We develop a novel subgraph-centric pre-training model, integrated with multi-view self-supervised learning, to capture urban generic knowledge from various modalities. Additionally, we propose two graph-based prompting methods—a manually-designed prompt and a task-learnable prompt—to inject explicit and implicit task-specific signals into the representations, thereby enhancing adaptability to diverse urban prediction tasks.
  • $\circ$
@@
    Extensive Empirical Validation: We demonstrate the effectiveness of GURPP through extensive experiments on six urban region prediction tasks across two cities. Our results show significant performance improvements, with an average increase of 13.57% on MAE and 16.24% on RMSE, highlighting the model’s strong capabilities in real-world urban prediction tasks.
Figure 3. Overview of GURPP framework. (1) Construct the urban region graph that integrates specific urban entities and semantic relationships. (2) Extract a subgraph of each region according to the graph pattern. (3) Learn the subgraph-enhanced region representations with multi-view fusion. (4) Adapt the region embeddings with a manually-designed prompt method and a task-learnable prompt learning method.
2. GURPP Framework
2.1. Preliminaries
@@
Once the mapping function $\zeta(\cdot)$ has been learned, it can be applied to various downstream prediction tasks, such as crime and check-in predictions. Let the task dataset be $\{(r_{i},y_{i})\}_{i=1}^{|\mathcal{R}|}$, where $y_{i}\in\mathbb{R}$ is a numerical indicator for region $r_{i}$. The downstream prediction tasks aim to predict $y_{i}$ based on the region embedding $\zeta(r_{i})$ using a simple regression model like a Ridge Regression (Hoerl and Kennard, 1970). However, learning effective region representations for specific tasks is challenging. First, task datasets are often very limited. This means the function $\zeta(\cdot)$ must work well in few-shot or even zero-shot situations. Second, urban spatial entities are diverse and large in number. This requires $\zeta(\cdot)$ to pull out general knowledge while still focusing on features specific to the task.
2.2. Graph-based pre-training and prompting
We propose GURPP, a unified framework that integrates urban graph pre-training with prompt learning. We formalize the GURPP framework in Eq. (1), where GURPP organizes spatial entities and inter-regional relationships within regions using a graph construction function $G(\cdot)$: During the pre-training stage $\mathcal{M}^{G}_{pre}$, the graph structure enables fine-grained feature extraction, such as capturing granular details like store brands within regions; In the prompting phase $\mathcal{M}^{G}_{pmt}$, GURPP dynamically assembles spatial entities and relationships for a region $r$ via graph prompts $\mathcal{P}$ to generate task-specific representations.
(1) $\zeta(\mathcal{P},r;\mathcal{R})=\mathcal{M}^{G}_{pre}\left(\mathcal{M}^{G}_{pmt}\left(\mathcal{P},G(r)\right),G(\mathcal{R})\right)$
GURPP integrates generalizable and task-specific knowledge within a unified framework. Its effectiveness can be demonstrated by recent advancements in graph prompt learning  (Wang et al., 2025; Yu et al., 2024b; Li et al., 2024b). Moreover, GURPP offers flexibility by supporting versatile forms of prompting for the same pre-trained model, such as manually-designed prompting (e.g., graph patterns  (Dong et al., 2017; Huang and Zitnik, 2020) and graph rules  (Fan et al., 2022a, b)) or task-learnable prompting (e.g.,  soft prompting  (Liu et al., 2024; Chen et al., 2023)), as long as the prompt operation conforms to the urban graph schema.
In the following sections, we detail the GURPP framework with a four-step workflow in Figure 3. First, an urban graph is built to unify heterogeneous spatial entities (Section 3.1). Second, region-specific subgraphs are extracted using structural and interaction-based patterns to identify region-specific patterns and relationships within urban data (Section 3.2). Third, a pre-trained multi-view learning model refines these subgraphs into generalizable embeddings (Section 3.3). Finally, task-specific knowledge is injected via prompt-tuning strategies (Section 4), enabling efficient adaptation without compromising the pre-trained backbone.
@@ 3.1. Urban Region Graph Construction
An urban region graph models the spatial structures and functional layouts of urban regions. Given a set of regions $\mathcal{R}$, the urban region graph, denoted as $G(\mathcal{R})=(V,E,T_{V},T_{E})$ (we simplify $G(\mathcal{R})$ to $G$ in the following), is a heterogeneous graph that captures multiple types of nodes and relationships. Here, $v\in V$ and $e\in E$ represent a node and an edge, respectively, while $\phi(v)\in T_{V}$ and $\psi(e)\in T_{E}$ are type functions that assign a specific type to a node or edge.
The urban region graph follows a schema that defines the relationships among various types of nodes and edges (see in Appendix A.1). Without loss of generality, we consider three types of spatial entities: POIs, roads, and junctions. Each spatial entity carries rich semantics; for instance, POIs have categories like shops and parks, and brands such as Dunkin’ and Starbucks. Roads include types such as highways and streets and junctions cover categories like roundabouts. This design effectively captures spatial structures and functional layouts at a granular level. For instance, two nearby regions containing coffee shops with brands of Blue Bottle and Dunkin’ are treated the same in existing statistic-based methods due to identical POI categories. In contrast, our region graph leverages semantic relations like BrandOf to distinguish them, enabling fine-grain region understanding, such as identifying high-end vs. casual commercial areas.
We construct the urban region graph from POI datasets and online maps. Since the graph contains heterogeneous nodes and edges, we first adopt the knowledge graph embedding technique TransR (Lin et al., 2015) to initialize the representations of all nodes and edges, aligning the heterogeneous information into a unified vector space.
@@ 3.2. Subgraph-centric Region Embedding
The region-induced subgraph is a subgraph of $G$ that includes the region node $r$ and matches the graph pattern $\mathcal{P}=(P_{V},P_{E},\theta)$. Here, $P_{V}\subseteq T_{V}$ and $P_{E}\subseteq T_{E}$ are sets of node and edge types, respectively, which filter entities and relationships specifically useful for region embeddings. The set $\theta\subseteq T_{V}$ indicates the node types that serve as termination points for subgraph extraction. As a result, the schema of the region-induced subgraph is a subset of the schema of the urban region graph. During the pre-training phase, we select all node and edge types and set $P_{V}=T_{V},P_{E}=T_{E}$, i.e., $\mathcal{P}=(T_{V},T_{E},\theta)$ with the aim of comprehensively characterising the rich information of a region and thereby learning general knowledge. The region-induced subgraph $g_{r}=(V_{r},E_{r},P_{V},P_{E})$ includes nodes $V_{r}$ and edges $E_{r}$ that can be reached from $r$ and whose types match $P_{V}$ and $P_{E}$. We extract $g_{r}$ recursively, starting with $V_{r}=\{r\}$ and $E_{r}=\varnothing$ and repeatedly expanding $V_{r}$ and $E_{r}$ until the termination points are reached. Here, the maximum boundary for subgraph extraction is set to the two-hop neighbors of the region node according to the graph schema.
@@
In Section 3.1, each node in the subgraph $g_{r}$ has been assigned a global initial feature vector $\vec{v}$ using TransR. We then encode each subgraph $g_{r}$ into a vector $\vec{h}_{r}$ to further incorporate the local features of region $r$. Specifically, we use a heterogeneous graph encoder $\textsf{gEnc}(\cdot)$ to extract the heterogeneous and generic patterns within $g_{r}$, ultimately outputting the contextualized embedding of the subgraph.
We choose the Heterogeneous Graph Transformer (HGT) (Hu et al., 2020) as the subgraph node encoder $\textsf{gEnc}(\cdot)$ due to its capability of capturing node and edge-type semantics. By using HGT, $\textsf{gEnc}(\cdot)$ characterizes the heterogeneity of input graph by maintaining a set of node and edge-type dependent parameters, and the output embedding of each node $\vec{v}^{HGT}$ that integrates the global initial embeddings generated by TransR, and node and edge-type semantics of the subgraph $g_{r}$. This trait aligns with our need to capture the semantic features of heterogeneous entities and relationships within each region-induced subgraph.
Obtaining the representations of each node in the subgraph, we further derive the graph-level representation of $g_{r}$ through two steps of aggregation. For each subgraph $g_{r}$, we first use $\text{SUM}(\cdot)$ function to aggregate the representations of nodes of the same type to integrate the impact of the number of entities on the representation, obtaining the node-type embedding $\vec{\tau}_{i}$: $\vec{\tau}_{i}=\sum_{\phi(v)=\tau_{i}}\vec{v}^{HGT}$, $\tau_{i}\in T_{V},\forall v\in V_{r}$. Then we concatenate all representations of node types, and encode them with a Linear layer to capture the interaction between various types of nodes: $\vec{h}_{r}=\text{Linear}\left(\text{CONCAT}_{|T_{V}|}(\vec{\tau}_{i})\right).$ The output $\vec{h}_{r}$ of $\textsf{gEnc}(\cdot)$ is the final embedding of region-induced subgraph $g_{r}$. Once the graph encoder is pretrained, any graph that conforms to the pattern $\mathcal{P}$ can be encoded by $\textsf{gEnc}(\cdot)$ to generate the corresponding representations, demonstrating high flexibility.
3.3. Multi-view Self-supervised Learning
We learn region representations in a self-supervised manner by integrating multiple views such as spatial proximity, region imagery and mobility patterns to characterize general region features.